Frida and Diego are both painters. At the time they met, Frida had sold 11 paintings and Diego had sold 37. From that point on, Frida sells 7 paintings per year and Diego sells 5 paintings per year. When they have sold the same number of paintings, how many paintings will each have sold?
Answer by richwmiller(17219)

f=11+7y
d=37+5y
11+7y=37+5y
2y=26
y=13
f=11+13*7=11+91=102
d=37+13*5=37+65=102
After 13 years they will each have sold 102 paintings.